Thermal Radiation: The Stefan-Boltzmann Law

Andy Chmilenko, 20310799 Instructor: Tan Dinh Section 1

(Dated: 2:30 pm Wednesday June 26, 2013)

I. PURPOSE

The purpose of this experiment is to verify the Stefan-Boltzmann Law, investigate aspects of blackbody radiation such as radiation rates from different surfaces and measure their emissivity (ε), as well as the absorption and transmission of thermal radiation as well as the inverse square law with respect to the radiative power of our Stefan-Boltzmann lamp.

II. ANALYSIS

A. Radiation rates from different Surfaces

Cube No. Thermistor Resistance Temperature  Sensor reading (mV)
(Ω)  (^{∘}K)  Black White Dull Al Polished Al
9  4692  388.7  17.4 16.0 2.6  0.9
8  5500  373.2  16.5 16.2 2.9  1.1
6  1391  345.6  8.9 9.0  2.6  0.9
5  2143  334.5  6.4 6.3  1.4  0.6

TABLE I: Thermistor readings along with the calculated Temperatures for several Leslie cubes, as well as the sensor readings from the Thermopile for all 4 sides of the cubes.

The distribution function as seen in Eq.1 for the wavelength dependence of emission of thermal radiation by a blackbody at temperature T using Planck’s theory of quantization

I(λ, T) = \frac{2πhc^{2}}{λ^{5} · e^{\frac{hc}{λkT}}}  (1)

and integrating this over all wavelengths the total power is simply defined as Eq.2

P = σT^{4}  (2)

Using the thermopile, several surfaces of different temperatures were measured, and the data was compiled into Table I.

III. CONCLUSION

Appendices

TABLE II: Resistance versus Temperature for the Thermal Radiation Cube

TABLE III: Resistance versus Temperature for the Thermal Radiation Cube